# ETGL-DDPG: A Deep Deterministic Policy Gradient Algorithm for Sparse Reward Continuous Control

### v1

@@ 6 Conclusions and Future Work
We have introduced the ETGL-DDPG algorithm with three components that improve the performance of DDPG for sparse-reward goal-conditioned environments. $\epsilon t$-greedy is a temporally-extended version of $\epsilon$-greedy using options generated by search. We prove that $\epsilon t$-greedy achieves a polynomial sample complexity under specific MDP structural assumptions. GDRB employs an extra buffer to separate successful episodes. The longest $n$-step return bootstraps from the Q-value of the final state in unsuccessful episodes and becomes a Monte Carlo update in successful episodes. ETGL-DDPG uses these components with DDPG and outperforms state-of-the-art methods, at the expense of about 1.5x wall-clock time w.r.t DDPG. The current limitation of our work is that we approximate visit counts through static hashing. For image-based problems such as real-world navigation, the future direction is to leverage dynamic hashing techniques such as normalizing flows (Papamakarios et al., 2021) as these tasks demand more intricate representation learning.
@@ A.1 $\epsilon t$-greedy Sample Efficiency : Proofs
We aim to show that the $\epsilon t$-greedy algorithm falls into the PAC-MDP category. According to Liu & Brunskill (2018), an algorithm $\mathcal{A}$ is PAC-MDP if the covering time induced by $\mathcal{A}$ is polynomially bounded. In Liu & Brunskill (2018), the authors further demonstrate that bounding the covering time requires bounding both the Laplacian eigenvalues and the stationary distribution over the states induced by the random walk policy. This is presented as Proposition A.1. According to Theorem 2, two conditions are satisfied: $N\leq\Theta(|\mathcal{S}||\mathcal{A}|)$ and a lower bound on the probability of the sampled option, $\mathcal{P}_{\mathcal{X}}\geq\frac{1}{\Theta(|\mathcal{S}||\mathcal{A}|)}$. These two conditions are necessary and are met by our problem setting and the exploration algorithm (Algorithm 1). To prove that $\mathcal{P}_{\mathcal{X}}\geq\frac{1}{\Theta(|\mathcal{S}||\mathcal{A}|)}$, we construct a worst-case tree structure $\mathcal{X}$, where we aim to identify the option induced by the tree $\mathcal{X}$ with the lowest probability, referred to informally as the “hardest option”. We then show that this lower bound satisfies the condition specified in Theorem 1.
@@
A.2 Exploration with a perfect model
@@
A.3 Implementation Details and Experimental Hyperparameters
Here, we describe the implementation details and hyperparameters for all methods used in this paper. All experiments were run on a system with 5 vCPU on a cluster of Intel Xeon E5-2650 v4 2.2GHz CPUs and one 2080Ti GPU. Table 3 displays the details for environments. Tables 2, 4, and 5 showcase the hyperparameters utilized in ETGL-DDPG and the baselines.
@@ A.3 Implementation Details and Experimental Hyperparameters
Table 4: Implementation details for SAC, TD3, and DDPG.
batch size 128 512
update frequency per step 12 2
action noise $\sim N(0,0.2)$ $\sim N(0,(0.3,0.05))$ $\sim N(0,(0.15))$
warmup period $2*10^{5}$ steps
replay buffer size $10^{6}$
learning rate $3*10^{-4}$
soft target updates $\tau$ $5*10^{-3}$
discount factor $\gamma$ 0.99
@@
A.4 ETGL-DDPG Algorithm
@@
A.5 Terminal states distribution
We analyze the order in which the agent visits different parts of the environment by examining the distribution of the last states in the episodes. To make it more visually appealing and easy to interpret, we only sample some of the episodes. The results for Wall-maze, U-maze, and Point-push can be found in Figures 7, 8, and 9, respectively. In Wall-maze, only $\epsilon t$-greedy and DOIE can effectively navigate to different regions of the environment and ultimately reach the goal area. Other methods often get trapped in one of the local optima and are unable to reach the goal. The reason some methods, such as TD3, have fewer points is that the agent spends a lot of time revisiting congested areas instead of exploring new ones. In U-maze, most methods can explore the majority of the environment. However, during the final stages of training, methods such as DDPG, SAC, and DDPG + intrinsic motivation have lower success rates and may end up in locations other than the goal areas. In Point-push, $\epsilon t$-greedy, $\epsilon z$-greedy, and DOIE first visit the lower section of the environment in the early stages. After that, they push aside the movable box and proceed to the upper section to visit the goal area. For the other methods, the pattern is almost the same, with occasional visits to the lower section.

### v2

@@ 6 Conclusions and Future Work
We have introduced the ETGL-DDPG algorithm with three orthogonal components that improve the performance of DDPG for sparse-reward goal-conditioned environments. $\epsilon t$-greedy is a temporally-extended version of $\epsilon$-greedy using options generated by search. We prove that $\epsilon t$-greedy achieves a polynomial sample complexity under specific MDP structural assumptions. GDRB employs an extra buffer to separate successful episodes. The longest $n$-step return bootstraps from the Q-value of the final state in unsuccessful episodes and becomes a Monte Carlo update in successful episodes. ETGL-DDPG uses these components with DDPG and outperforms state-of-the-art methods, at the expense of about 1.5x wall-clock time w.r.t DDPG. The current limitation of our work is that we approximate visit counts through static hashing. For image-based problems such as real-world navigation, the future direction is to leverage dynamic hashing techniques such as normalizing flows (Papamakarios et al., 2021) as these tasks demand more intricate representation learning. Additionally, $\epsilon t$-greedy and GDRB are designed for deterministic domains and would require adaptation for stochastic environments. Extending these components to stochastic domains represents another promising direction for future research.
@@ A.1 $\epsilon t$-greedy Sample Efficiency : Proofs
We aim to show that the $\epsilon t$-greedy algorithm falls into the PAC-MDP category. According to Liu & Brunskill (2018), an algorithm $\mathcal{A}$ is PAC-MDP if the covering time induced by $\mathcal{A}$ is polynomially bounded. In Liu & Brunskill (2018), the authors further demonstrate that bounding the covering time requires bounding both the Laplacian eigenvalues and the stationary distribution over the states induced by the random walk policy. This is presented as Proposition A.1. According to Theorem 2, two conditions are satisfied: $N\leq\Theta(|\mathcal{S}||\mathcal{A}|)$ and a lower bound on the probability of the sampled option, $\mathcal{P}_{\mathcal{X}}\geq\frac{1}{\Theta(|\mathcal{S}||\mathcal{A}|)}$. These two conditions are necessary and are met by our problem setting and the exploration algorithm (Algorithm 1). To prove that $\mathcal{P}_{\mathcal{X}}\geq\frac{1}{\Theta(|\mathcal{S}||\mathcal{A}|)}$, we construct a worst-case tree structure $\mathcal{X}$, where we aim to identify the option induced by the tree $\mathcal{X}$ with the lowest probability, referred to informally as the “hardest option". We then show that this lower bound satisfies the condition specified in Theorem 1.
@@
A.2 n-step Methods Comparison
In this section, we compare two update rules for Q-values in DDPG: the longest n-step return and the average return over 1 to 8 steps (avg8-step). The results are shown in Figure 9. Both methods performed similarly overall; however, in the Wall-maze environment, the longest n-step return outperformed the avg8-step approach. The avg8-step method converged more quickly in U-maze and Point-push due to the stabilizing effect of averaging multiple updates.
A.3 Comparison of GDRB and Hindsight Experience Replay (HER)
In this section, we compare two strategies for storing transitions in the replay buffer: GDRB and HER. Figure 8 shows the results in the navigation environments. In Wall-maze, HER achieves success rates above zero during the early stages of training; however, it fails to maintain any success in later stages. This observation is consistent with the results reported by Trott et al. (2019) in the same environment. In the other two environments, both methods exhibit comparable performance, but HER discovers paths to the goal more quickly, as it leverages reward shaping in unsuccessful episodes to provide additional information about the structure of the environment.
(a) Wall-maze
(b) U-maze
(c) Point-push
Figure 8: The comparison between two replay buffer strategies: GDRB and HER.
(a) Wall-maze
(b) U-maze
(c) Point-push
Figure 9: The comparison between two update rules for Q-values: the longest n-step return and the average return over 1 to 8 steps (avg8-step).
A.4 Exploration with a Perfect Model
@@
A.5 Implementation Details and Experimental Hyperparameters
Here, we describe the implementation details and hyperparameters for all methods used in this paper. All experiments were run on a system with 5 vCPU on a cluster of Intel Xeon E5-2650 v4 2.2GHz CPUs and one 2080Ti GPU. Table 2 displays the details for environments. Tables 1, 4, and 3 showcase the hyperparameters utilized in ETGL-DDPG and the baselines.
@@ A.5 Implementation Details and Experimental Hyperparameters
Table 4: Implementation details for SAC, TD3, and DDPG.
batch size 128 512
update frequency per step 12 2
action noise $\sim N(0,0.2)$ $\sim N(0,(0.3,0.05))$ $\sim N(0,(0.15))$
warmup period $2*10^{5}$ steps
replay buffer size $10^{6}$
learning rate $3*10^{-4}$
soft target updates $\tau$ $5*10^{-3}$
discount factor $\gamma$ 0.99
A.6 ETGL-DDPG Algorithm
@@
A.7 Analyzing the Impact of $N$ on $\epsilon t$-greedy and $\epsilon z$-greedy Exploration
The tree budget $N$ upper bounds the option length of $\epsilon t$-greedy due to the fact that the longest path between nodes in the tree is shorter or equal to the number of nodes in the tree. This is analogous to the role of $N$ in $\epsilon z$-greedy, where a uniform distribution $z(n)=\mathbbm{1}_{n\leq N}/N$ is used. To evaluate both methods, we assess environment coverage under varying budget sizes, calculating the coverage after 1 million training frames. Table 5 shows the results: $\epsilon t$-greedy consistently achieves greater coverage than $\epsilon z$-greedy across all environments and budget sizes. Additionally, $\epsilon t$-greedy demonstrates improved coverage as the budget increases. In contrast, increasing the budget for $\epsilon z$-greedy does not consistently improve coverage and can even decrease it in some cases. This highlights the advantages of directed exploration over undirected methods, particularly in complex environments with numerous obstacles, such as Wall-maze.
budget $N$
$\epsilon z$-greedy $\epsilon t$-greedy
Wall-maze U-maze Point-push Wall-maze U-maze Point-push
5 0.36 0.55 0.36 0.76 0.94 0.40
10 0.38 0.91 0.38 0.97 0.91 0.41
15 0.34 0.85 0.39 0.65 0.94 0.42
20 0.30 0.84 0.40 0.83 0.94 0.48
25 0.28 0.86 0.40 1 0.95 0.47
30 0.27 0.83 0.39 1 0.97 0.51
35 0.25 0.82 0.40 1 0.95 0.53
40 0.24 0.82 0.40 1 0.97 0.55
45 0.22 0.85 0.41 1 0.96 0.64
50 0.22 0.79 0.40 1 0.97 0.73
A.8 Terminal States Distribution
We analyze the order in which the agent visits different parts of the environment by examining the distribution of the last states in the episodes. To make it more visually appealing and easy to interpret, we only sample some of the episodes. The results for Wall-maze, U-maze, and Point-push can be found in Figures 11, 12, and 13, respectively. In Wall-maze, only $\epsilon t$-greedy and DOIE can effectively navigate to different regions of the environment and ultimately reach the goal area. Other methods often get trapped in one of the local optima and are unable to reach the goal. The reason some methods, such as TD3, have fewer points is that the agent spends a lot of time revisiting congested areas instead of exploring new ones. In U-maze, most methods can explore the majority of the environment. However, during the final stages of training, methods such as DDPG, SAC, and DDPG + intrinsic motivation have lower success rates and may end up in locations other than the goal areas. In Point-push, $\epsilon t$-greedy, $\epsilon z$-greedy, and DOIE first visit the lower section of the environment in the early stages. After that, they push aside the movable box and proceed to the upper section to visit the goal area. For the other methods, the pattern is almost the same, with occasional visits to the lower section.